OpenStudy (anonymous):

simplifty: log3(y^7) = 14

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

cold you please provide me with the steps

OpenStudy (anonymous):

oops!sry buddy....so yea, take 3 power on both sides, so u get...y^7=3^14 so y=3^2=9. get it?

OpenStudy (anonymous):

take 3 power on both sides? did you mean to divide by sides by 7, so we get 3^2

OpenStudy (anonymous):

\[\text{Log}\left[3, y^7\right]=14, y^7=3^{14},\sqrt[7]{y{}^{\wedge}7}=\sqrt[7]{3^{14}},y^{\left(7*\frac{1}{7}\right)}=3^{\left(14*\frac{1}{7}\right)},y=3^2,y=9 \]